OpenStudy (anonymous):

Solve the equation without a calculator: (logx)^2=log(x^2)

OpenStudy (callisto):

(logx)^2=log(x^2) (logx)^2=2logx (logx)^2-2logx = 0 logx (logx - 2) =0 logx =0 or logx -2 = 0 <- solve these 2 equations.

OpenStudy (anonymous):

When it's solved it's like... why didn't I think of that, sigh.... Thanks @Callisto

OpenStudy (callisto):

Welcome. It's just a matter of experience.

OpenStudy (anonymous):

\[\log(x) \times \cancel{\log(x)} = 2 \cancel {\log(x)}\] logx = 2

OpenStudy (anonymous):

x=1 and x=100

OpenStudy (anonymous):

Yes you are right..

OpenStudy (callisto):

@waterineyes Honestly, you shouldn't cross out the variable like that, even though you can do that by putting it = 0.

OpenStudy (anonymous):

And I guess... in reference to the experience

OpenStudy (anonymous):

Ya I got it Callisto because the equation is quadratic and you will get two values but I am getting one..

OpenStudy (callisto):

@purplec16 lol It's just a trick I used in the exam

OpenStudy (callisto):

But I don't recommend you do that. Just stick to the way I posted. It's safer.

OpenStudy (anonymous):

Oki, thanks once again!

OpenStudy (callisto):

x^2 = 2x x^2 - 2x =0 x( x-2) =0 x=0 or x-2 =0 ^^ Use this method!! But from that you can see that \(x^2 = 2x\) \(x \times x = 2x \) ''Cancel'' the x \(x=2\) or \(x=0\)

OpenStudy (anonymous):

what happens when you take the log of a log like log(logx)

OpenStudy (callisto):

The value would become smaller?!

OpenStudy (anonymous):

No I mean like log(logx)=2...

OpenStudy (anonymous):

10^100=x?

OpenStudy (anonymous):

or is this a cancellation?

OpenStudy (callisto):

log(logx)=2 logx = 10^2 logx = 100 x = 10^100

OpenStudy (anonymous):

Yay me! I'm getting it! Thanks for your help! Your the bestest!

OpenStudy (callisto):

Welcome. I'm not the best though

OpenStudy (anonymous):

Yeah you are!